OpenStudy (anonymous):

8x – 12 = 4x + 24 I Know that the answer is x=9 but can you guys help me show the work. It is hard for me to put it into words.

OpenStudy (waheguru):

Do you need help in putting the answer in words ot numners?

OpenStudy (anonymous):

into words and numbers like this is how I did my other one Subtract 69 from both sides. then you get: -24 = 3b Divide each side by 3. You get b= -8 and that is the answer

OpenStudy (anonymous):

8x-12 = 4x+24 Add twelve to both sides: 8x-12+12=4x+24+12, 8x=4x+36 Subtract 4x from both sides: 8x-4x=4x-4x+36, 4x=36 Divide both sides by four: 4x/4=36/4, x=9

OpenStudy (anonymous):

Thanks so much that is what I was trying to write it just would not come ot

OpenStudy (waheguru):

np